It was established by the Biology Department years ago that the
squirrels on campus are quite healthy, and average 4 pounds each. However, the
squirrels this year have seemed even healthier, so the Bio Dept decided to test them to
see if their average weight had increased. A recent sample of 16 squirrels was found to
have a mean weight of 4.3 pounds, with a standard deviation of 1.2 pounds. Do the
test for the Bio Dept.
Answer #1

x̅ = 4.3, s = 1.2, n = 16  

Null and Alternative hypothesis:  

Ho : µ = 4

H1 : µ = 4

Test statistic:  

t = (x̅- µ)/(s/√n) = (4.3 - 4)/(1.2/√16) = 1.0

df = n-1 = 15

p-value = T.DIST.RT(1, 15) = 0.1666

Decision:  

p-value > 0.05, Do not reject the null hypothesis  

Conclusion:  

There is not enough evidence to conclude that the average weight had increased.